ganeshie8 (ganeshie8):

show that \[\sum\limits_{d\mid n}d\tau(d) = \sum\limits_{d\mid n}(n/d)\sigma(d)\]

ganeshie8 (ganeshie8):

it is given, \(n\) is of form \(p^k\) where \(p\) is a prime

ganeshie8 (ganeshie8):

in other words \(n\) is power of a prime

OpenStudy (pawanyadav):

As always ...explain the term first pls So that we can also understand it

ganeshie8 (ganeshie8):

Here is the idenity incase if the latex has trouble displaying in main question \[\sum\limits_{d\mid n}d\tau(d) = \sum\limits_{d\mid n}(n/d)\sigma(d)\]

ganeshie8 (ganeshie8):

\(\tau(n)\) gives the "number" of positive divisors of \(n\) \(\sigma(n)\) gives the "sum" of positive divisors of \(n\)

ganeshie8 (ganeshie8):

the sum \(\sum\limits_{d\mid n}f(d)\) means : "Sum the values f(d) as d runs over all the positive divisors of n"

ganeshie8 (ganeshie8):

example : \[\sum\limits_{d\mid 10}f(d) = f(1)+f(2)+f(5)+f(10)\] because positive divisors of 10 are {1, 2, 5, 10}

OpenStudy (pawanyadav):

Can I say LHS as f(1)+ f(p)+ f(p^2)+f(p^3)+...............+f(p^n)

OpenStudy (pawanyadav):

f(1)+f(p)+f(p^2)+...........+f(p^k)

ganeshie8 (ganeshie8):

LHS is \(\sum\limits_{d\mid n}d\tau(d)\) i think we get : \[1\tau(1) + p\tau(p) + p^2\tau(p^2) + \cdots + p^k\tau(p^k)\]

OpenStudy (pawanyadav):

Yes, my i understand my mistake

OpenStudy (mayankdevnani):

@ganeshie8

ganeshie8 (ganeshie8):

yes i think we may use those results...

OpenStudy (pawanyadav):

So LHS would be 1+2p+3(p^2)+.........+(k+1)(p^k)

OpenStudy (mayankdevnani):

happy to help you :)

ganeshie8 (ganeshie8):

Oh nice, i think that is correct !

OpenStudy (pawanyadav):

But T(1)=1 T(p)=2 As p a prime it have two +ve divisor

OpenStudy (pawanyadav):

Yes but what next

ganeshie8 (ganeshie8):

So, according to you : \(\sum\limits_{d\mid n}d\tau(d)=1\tau(1) + p\tau(p) + p^2\tau(p^2) + \cdots + p^k\tau(p^k)\\~\\~\\ =1*1+p*2+p^2*3+\cdots + p^k(k+1)\\~\\ =1+2p+3p^2+\cdots+(k+1)p^k \)

ganeshie8 (ganeshie8):

I don't know whats next, still thinking...

OpenStudy (pawanyadav):

Is it a aritho-geo-progression

OpenStudy (pawanyadav):

Both AP and GP

ganeshie8 (ganeshie8):

yeah it may be sufficient if we could show that RHS also equals this aritho-geo-progression

OpenStudy (pawanyadav):

I can solve it ...OK now starts with RHS

OpenStudy (pawanyadav):

So RHS = n S(1) + n/p S(p)+.......+n/n S(n)

OpenStudy (pawanyadav):

Where S represents Sigma And n=p^k

ganeshie8 (ganeshie8):

yes \(\sigma(p^k) = 1+p+p^2+\cdots+p^k\)

ganeshie8 (ganeshie8):

because the positive divisors of \(p^k\) are simply \(\{1,p,p^2,\ldots,p^k\}\)

ganeshie8 (ganeshie8):

That looks nice! the sum should be from i=0 to i=k right ?

imqwerty (imqwerty):

oops sooryy yea that should be i=k :)

ganeshie8 (ganeshie8):

\(\tau(p^k) = k+1\)

OpenStudy (pawanyadav):

How to take summation of that that series.

imqwerty (imqwerty):

ok i'll edit the solution edited solution-> my work so far- \(\sum\limits_{d\mid n}d\tau(d) = \sum\limits_{d\mid n}(n/d)\sigma(d)\) We know that \(n=p^k\) divisors of n will be->\(1, p, p^2, p^3...p^{k}\) putting these values in LHS we get a series like this-> \[1 \tau(1)+p \tau(p) + p^2 \tau(p^2)....p^{k} \tau(p^{k}) \] \[=>1 +2p+3p^2 +4p^3......(k+2)p^{k+1}\]\[=> \sum_{i=0}^{k}(i+1)p^i\] now the RHS \(\sum\limits_{d\mid n}(n/d)\sigma(d)\) \(\sum\limits_{d\mid n}(p^k /d)\sigma(d)\) putting d=\(1,~p,~p^2,~p^3...p^{k+1}\) we get a series like this-> \(p^k,~p^{k-1}[1+p], ~p^{k-2}[1+p+p^2],.....p^0[1+p+p^2..p^k]\) opening the brackets the series looks like this-> \(p^k,~p^{k}+p^{k-1}, ~ p^k+p^{k-1}+p^{k-2},.......,p^k+p^{k-1}..+p^{k-k}\) taking the summation we will get p^k we will get it k+1 times so we can write the RHS as\[\sum_{i=0}^{k}(i+1)p^i\] \(LHS=RHS\) hence proved

OpenStudy (pawanyadav):

Can you tell me pls how to take summation of that series?

imqwerty (imqwerty):

which series @pawanyadav ?

OpenStudy (pawanyadav):

That we found in RHS